OpenStudy (anonymous):

Evaluate: If a = 42 and b = 2, then find ab.

OpenStudy (alexandervonhumboldt2):

a*b=substitute=42*2=solve

OpenStudy (anonymous):

84 40 21 20

OpenStudy (alexandervonhumboldt2):

what is 42*2 @Tankkerking ?

OpenStudy (anonymous):

84

OpenStudy (alexandervonhumboldt2):

correct thus your answer is 84

OpenStudy (anonymous):

whats this one is ahrder

OpenStudy (alexandervonhumboldt2):

substitute 77 for F

OpenStudy (alexandervonhumboldt2):

c=5/9*(77-32)=5/9*45=5*5=25

OpenStudy (anonymous):

ok

OpenStudy (alexandervonhumboldt2):

substitute: 4.25*16+15/100*250=68+37.5=solve

OpenStudy (alexandervonhumboldt2):

65*7=multiply and get your answer